### v2

@@ Abstract
We study the differential geometry of the fixed-rank core covariance manifold. According to Hoff, McCormack, and Zhang [J. R. Stat. Soc., B: Stat., 85 (2023), pp. 1659–1679], every covariance matrix $\Sigma$ of $p_{1}\times p_{2}$ matrix-variate data uniquely decomposes into a separable component $K$ and a core component $C$. Such a decomposition also exists for rank-$r$ $\Sigma$ if $p_{1}/p_{2}+p_{2}/p_{1}<r$, with $C$ sharing the same rank. They posed an open question on whether a partial-isotropy structure can be imposed on $C$ for high-dimensional covariance estimation. We address this question by showing that a partial-isotropy rank-$r$ core is a non-trivial convex combination of a rank-$r$ core and $I_{p}$ for $p:=p_{1}p_{2}$, motivating the study of rank-$r$ cores. For fixed $r>p_{1}/p_{2}+p_{2}/p_{1}$, we prove that the set of rank-$r$ cores, $\mathcal{C}_{p_{1},p_{2},r}^{+}$, is a compact, smooth, embedded submanifold of the set of rank-$r$ positive semi-definite matrices, except for a measure-zero subset associated with canonical decomposability. When $r=p$, the set of full-rank cores $\mathcal{C}_{p_{1},p_{2}}^{++}$ is itself a smooth manifold. Moreover, the positive definite cone $\mathcal{S}_{p}^{++}$ is diffeomorphic to the product of the Kronecker and core covariance manifolds, providing new geometric insight into $\mathcal{S}_{p}^{++}$ via separability. Differential geometric quantities, such as the differential of the diffeomorphism, as well as the Riemannian gradient and Hessian operator on $\mathcal{C}_{p_{1},p_{2}}^{++}$ and the manifolds used in constructing $\mathcal{C}_{p_{1},p_{2},r}^{+}$, are also derived. Lastly, we propose a partial-isotropy core shrinkage estimator for matrix-variate data, supported by numerical illustrations.
@@
canonical decomposability, core covariance matrix, fixed-rank, Kronecker-core decomposition, Riemannian manifold optimization, separable covariance matrix.
{AMS}
14P05, 15B99, 53C55, 58C15, 62H12.
@@ 1 Introduction
Symmetric positive semi-definite (PSD) matrices arise in a wide range of modern applications. For example, many non-Euclidean data are often represented as PSD matrices, e.g., brain connectivity analysis [huettel2014], diffusion tensor imaging [bihan1991, dryden2009], and tomography [rau2018]. In statistics, PSD matrices commonly appear as covariance matrices, typically assumed to be strictly positive definite (PD). In the analysis of such data or the covariance estimation using the Riemannian geometry of the PD cone or its submanifolds, e.g., [bouchard2021, han2021, wiesel2012, mccormack2025, musolas2021], the Euclidean metric is not suitable as geodesics leave the space in finite time, resulting in non-PSD interpolations. Therefore, various metrics have been proposed for the PD cone, including affine-invariant [skovgaard1984, pennec2006, moakher2005], log-Euclidean [arsigny2007], log-Cholesky [lin2019], Bures-Wasserstein [bhatia2019, asuka2011, dowson1982], and a product metric with one metric on positive diagonal matrices and one metric on full-rank correlation matrices [thanwerdas2022]. While these metrics are defined on the PD cone, the quotient geometry has been studied for fixed or bounded rank PSD matrices [thanwerdas2023] and correlation matrices [chen2025].
In covariance estimation based on the Riemannian frameworks, the choice of the parameter space and metric depends on the assumed covariance model and the data type. As an example, for $p$-dimensional vector data, the parameter space of the population covariance matrix $\Sigma$ is typically considered as the PD cone of the order $p$ [musolas2021, han2021], denoted $\mathcal{S}_{p}^{++}$, where any aforementioned metric for the PD cone can be adopted. On the other hand, for $p_{1}\times p_{2}$ matrix-variate data, e.g., microarray data [allen2012], phonetic data [pigoli2014], and audio data [warden2018], a separable (Kronecker) covariance model [dawid1981] is commonly used. Namely, for a zero-mean $p_{1}\times p_{2}$ random matrix $Y$, its $p_{1}p_{2}\times p_{1}p_{2}$ covariance matrix $\Sigma$ is formulated as
(1) $\Sigma=V[Y]\equiv V[\text{vec}(Y)]=\Sigma_{2}\otimes\Sigma_{1},$
where $\Sigma_{1}\in\mathcal{S}_{p_{1}}^{++}$ and $\Sigma_{2}\in\mathcal{S}_{p_{2}}^{++}$ correspond to row and column covariance matrices, respectively. Here $\otimes$ denotes the Kronecker product. Note that we assume $p_{1},p_{2}\geq 2$ to emphasize the matrix structure in this article. It follows that
@@
enabling the separate inference of correlation structures of row and column factors [dawid1981, wang2022]. Under this model, there are at most $O(p_{1}^{2}+p_{2}^{2})$ correlations between variables, thereby making $\Sigma$ parsimonious. We denote the space of such separable covariance matrices by $\mathcal{S}_{p_{1},p_{2}}^{++}$. As a submanifold of $\mathcal{S}_{p}^{++}$ for $p:=p_{1}p_{2}$, [simonis2025, mccormack2025] proposed the estimation of the separable covariance matrix under the affine-invariant geometry.
@@
However, as $p$ grows, the separability assumption on $\Sigma$ as in (1) may oversimplify its correlation structure, allowing at most $O(p_{1}^{2}+p_{2}^{2})=o(p^{2})$ correlations, whereas $O(p^{2})$ correlations for the unstructured $\Sigma$. Hence, the separability assumption is often inappropriate, as also pointed out by [gneiting2002, gneiting2007]. As a departure from this assumption, [hoff2023] introduced the core covariance matrix. They showed that every $\Sigma\in\mathcal{S}_{p}^{++}$ admits a unique decomposition into a separable component $K$, representing the most separable part of $\Sigma$, and a core component $C$, whitened $\Sigma$ via the identifiable square root $K^{1/2}$ of $K$, e.g., symmetric square root and Cholesky factor. Namely, $\Sigma=K^{1/2}CK^{1/2,\top}$, referred to as a Kronecker-core decomposition (KCD). Also, $\Sigma\in\mathcal{S}_{p_{1},p_{2}}^{++}$ if and only if $C=I_{p}$. Such a decomposition may also exist for rankr-r $\Sigma$ if $p_{1}/p_{2}+p_{2}/p_{1}<r$ [drton2021, derksen2021, soloveychik2016], with $C$ sharing the same rank as $K^{1/2}$ is non-singular. By Proposition $5$ of [hoff2023], the dimension of the space of full-rank $C$ is $O(p^{2})$, whereas that of $\mathcal{S}_{p_{1},p_{2}}^{++}$ is $O(p_{1}^{2}+p_{2}^{2})=o(p^{2})$. Thus, in a high-dimensional regime where the sample size $n$ is smaller than the dimension of variables $p$, the estimation of $\Sigma$ may be either numerically or statistically unstable without any structural assumption on $C$.
@@
As discussed by [hoff2023], one remedy is to introduce a partial-isotropy rankr-r structure to $C$, which commonly arises in factor analysis [basilevsky1994, bartholomew2011]. Specifically, $\Omega\in\mathcal{S}_{p}^{++}$ has such a structure if $\lambda_{1}(\Omega)\geq\cdots\geq\lambda_{r}(\Omega)>\lambda_{r+1}(\Omega)=\cdots=\lambda_{p}(\Omega)>0$. Such $\Omega$ can be equivalently formulated as $AA^{\top}+cI_{p}$ for some $A\in\real^{p\times r}$ of full-column rank and constant $c>0$. Nevertheless, [hoff2023] did not pursue the estimation using this structure themselves, and left it as an open question, as characterizing such a core is crucial. In this article, we shall answer this question. It turns out that if $C$ exhibits a partial-isotropy rankr-r structure for fixed $r>p_{1}/p_{2}+p_{2}/p_{1}$, $C$ is a non-trivial convex combination of a rankr-r core and a trivial core $I_{p}$. As discussed in Section 2.1 and 6, the coefficient on $I_{p}$ quantifies how far $\Sigma$ is from being separable, improving the interpretability compared to an unstructured $C$. To incorporate a partial-isotropy $C$ into the estimation of $\Sigma$, we need a proper understanding of the space of rankr-r cores, denoted $\mathcal{C}_{p_{1},p_{2},r}^{+}$, motivating the study of its geometry.
@@
Therefore, this article is devoted to establishing the geometry of $\mathcal{C}_{p_{1},p_{2},r}^{+}$. Our contributions are summarized in three main strands. First, we show that $\mathcal{C}_{p_{1},p_{2},r}^{+}$ is a compact, smooth, embedded submanifold of $\mathcal{S}_{p,r}^{+}$, the set of rankr-r PSD matrices. A key observation for the proof is that if $C=AA^{\top}\in\mathcal{C}_{p_{1},p_{2},r}^{+}$ for $A=[\text{vec}(A_{i}),\ldots,\text{vec}(A_{r})]$ with $A_{i}\in\real^{p_{1}\times p_{2}}$, Proposition $3$ of [hoff2023] implies that
@@
Define the diffeomorphism $\varphi_{p_{1},p_{2},r}$ that maps $(A_{1},\ldots,A_{r})\in(\real^{p_{1}\times p_{2}})^{r}$ to $a_{1},\ldots,a_{r}]\in\real^{p\times r}$ for $a_{i}=\text{vec}(A_{i})$. The set $\mathcal{D}_{p_{1},p_{2},r}$ of $(A_{1},\ldots,A_{r})$ satisfying the above is shown to be a submanifold of the smooth manifold $(\real^{p_{1}\times p_{2}})_{*}^{r}$, the set of $X=(X_{1},\ldots,X_{r})\in(\real^{p_{1}\times p_{2}})^{r}$ such that $\varphi_{p_{1},p_{2},r}(X)$ is of rankr-r. The proof is straightforward when $r=p$, but the rank-deficient case requires additional technical work. Namely, we have to remove the closed measure-zero set $\mathcal{V}_{p_{1},p_{2},r}$ consisting of canonically decomposable matrices. The canonical decomposability notion arises in the study of the threshold on $r$ for which a generic $\Omega\in\mathcal{S}_{p,r}^{+}$ admits the KCD [derksen2021, soloveychik2016]. Here $(A_{1},\ldots,A_{r})$ is canonically decomposable if there exist non-singular matrices $(P,Q)$ such that each $PA_{i}Q^{-1}$ has a non-trivial block-diagonal form. We illustrate in Section 4.1 how $\mathcal{V}_{p_{1},p_{2},r}$ prevents $\mathcal{D}_{p_{1},p_{2},r}$ from being a smooth manifold and is a closed set of measure zero. This issue does not arise when $r=p$ as shown in Section 3. After removing this set, $\mathcal{D}_{p_{1},p_{2},r}$ is shown to be a smooth manifold. Observing further that $\mathcal{C}_{p_{1},p_{2},r}\equiv\varphi_{p_{1},p_{2},r}(\mathcal{D}_{p_{1},p_{2},r})$ is embedded in the set of rankr-r $p\times r$ matrices, that the orthogonal group $\mathcal{O}_{r}$ acts smoothly and freely on $\mathcal{C}_{p_{1},p_{2},r}$, and that $\mathcal{C}_{p_{1},p_{2},r}/\mathcal{O}_{r}$ is diffeomorphic to $\mathcal{C}_{p_{1},p_{2},r}^{+}$, we show that $\mathcal{C}_{p_{1},p_{2},r}^{+}$ is a smooth manifold and identify its tangent space. The proof strategy is outlined more clearly with the case $r=p$, where $\mathcal{C}_{p_{1},p_{2}}^{++}\equiv\mathcal{C}_{p_{1},p_{2},p}^{+}$, in Section 3, and then extended to the case where $r<p$ in Section 4.2.
@@
The next contribution is to study the differential geometry of $\mathcal{C}_{p_{1},p_{2}}^{++}$, $\mathcal{C}_{p_{1},p_{2},r}$, and the quotient manifold $\mathcal{C}_{p_{1},p_{2},r}/\mathcal{O}_{r}$. Let $k(\Sigma)$ and $c(\Sigma)$ denote the separable and core components of $\Sigma$, respectively. We refer to $k$ and $c$ as the Kronecker and core maps defined on $\mathcal{S}_{p}^{++}$, respectively. With the map $f:\Sigma\in\mathcal{S}_{p}^{++}\rightarrow(k(\Sigma),c(\Sigma))\in\mathcal{S}_{p_{1},p_{2}}^{++}\times\mathcal{C}_{p_{1},p_{2}}^{++}$, we show that $\mathcal{S}_{p}^{++}$ is diffeomorphic to $\mathcal{S}_{p_{1},p_{2}}^{++}\times\mathcal{C}_{p_{1},p_{2}}^{++}$ via the map $f$ in Section 5.1. Therefore, we provide a new insight into the smooth structure of $\mathcal{S}_{p}^{++}$ via separability. We also compute the differentials of $f$ and its inverse $f^{-1}$. Under the Euclidean metric, we derive the Riemannian gradient and Hessian operator on $\mathcal{C}_{p_{1},p_{2}}^{++}$ in Section 5.2. The same is done for $\mathcal{C}_{p_{1},p_{2},r}$ under the same metric, and for the quotient manifold $\mathcal{C}_{p_{1},p_{2},r}/\mathcal{O}_{r}$ under the induced quotient metric in Section 5.3.
@@
Finally, using the geometry of $\mathcal{C}_{p_{1},p_{2},r}$, we propose a partial-isotropy core shrinkage estimator (PICSE) for the true covariance $\Sigma$ with a partial-isotropy core in Section 6. Namely, $\Sigma=K^{1/2}((1-\lambda)AA^{\top}+\lambda I_{p})K^{1/2,\top}$, where $K=K^{1/2}K^{1/2,\top}\in\mathcal{S}_{p_{1},p_{2}}^{++}$, $\lambda\in(0,1)$ and $A\in\mathcal{C}_{p_{1},p_{2},r}$. This answers the open question posed by [hoff2023] on how a partial-isotropy core can be incorporated into estimating $\Sigma$. We propose an alternating minimization algorithm to compute PICSE. In updating $A$, we leverage the curvature of the objective function on $\mathcal{C}_{p_{1},p_{2},r}$ via second-order Riemannian manifold optimization using the results in Section 5.3, with some suitable retraction. In Section 7, we numerically illustrate that PICSE outperforms existing covariance estimators for matrix-variate data, and some baseline methods.
@@
The rest of the article is organized as follows. Section 1.1 introduces notations used throughout this article. In Section 2, we review some preliminaries, including the KCD, Riemannian manifolds, quotient manifolds, and algebraic geometry. In Section 3, we prove that $\mathcal{C}_{p_{1},p_{2}}^{++}$ is a compact, smooth, embedded submanifold of $\mathcal{S}_{p}^{++}$. When $p_{1}/p_{2}+p_{2}/p_{1}<r<p$, it is shown that $\mathcal{C}_{p_{1},p_{2},r}^{+}$ is a compact, smooth, embedded submanifold of $\mathcal{S}_{p,r}^{+}$ in Section 4 after removing the set of canonically decomposable matrices, using the proof strategy developed in Section 3. In Section 5, we establish the diffeomorphic relationship between $\mathcal{S}_{p}^{++}$ and $\mathcal{S}_{p_{1},p_{2}}^{++}\times\mathcal{C}_{p_{1},p_{2}}^{++}$. We also derive the differential geometric quantities relevant to $\mathcal{C}_{p_{1},p_{2}}^{++}$, $\mathcal{C}_{p_{1},p_{2},r}$, and $\mathcal{C}_{p_{1},p_{2},r}/\mathcal{O}_{r}$ under the Euclidean metric. Leveraging the geometry of $\mathcal{C}_{p_{1},p_{2},r}$, the partial isotropy core shrinkage estimator (PICSE) is proposed in Section 6, supported by numerical illustrations in Section 7. Section 8 concludes the article with a discussion. All the omitted proofs are deferred to Appendix A. The formulas of Euclidean derivative and Hessian operator associated with computing PICSE are provided in Appendix B. The additional figures that illustrate the results in Section 7 are given in Appendix C.
1.1 Notation

### v3

@@ Abstract
We study the geometry of the fixed-rank core covariance manifold and propose a novel covariance estimator for matrix-variate data leveraging this geometry. To generalize the separable covariance model, Hoff, McCormack, and Zhang (2023) showed that every covariance matrix $\Sigma$ of $p_{1}\times p_{2}$ matrix-variate data uniquely decomposes into a separable component $K$ and a core component $C$. Such a decomposition also exists for rank-$r$ $\Sigma$ if $p_{1}/p_{2}+p_{2}/p_{1}<r$, with $C$ sharing the same rank. They posed an open question on whether a partial-isotropy structure can be imposed on $C$ for high-dimensional covariance estimation. We address this question by showing that a partial-isotropy rank-$r$ core is a non-trivial convex combination of a rank-$r$ core and $I_{p}$ for $p:=p_{1}p_{2}$. This motivates studying the geometry of the space of rank-$r$ cores, $\mathcal{C}_{p_{1},p_{2},r}^{+}$. We show that $\mathcal{C}_{p_{1},p_{2},r}^{+}$ is a smooth manifold, except for a measure-zero subset, whereas $\mathcal{C}_{p_{1},p_{2}}^{++}:=\mathcal{C}_{p_{1},p_{2},p}^{+}$ is itself a smooth manifold. The geometric properties, including smoothness of the positive definite cone via separability and the Riemannian gradient and Hessian operator relevant to $\mathcal{C}_{p_{1},p_{2},r}^{+}$, are also derived. Using this geometry, we propose a partial-isotropy core shrinkage estimator for matrix-variate data, supported by numerical illustrations.
separable covariance matrix,
@@
[class=MSC]
keywords:
[…]
@@ 1 Introduction
Symmetric positive semi-definite (PSD) matrices arise in a wide range of modern applications. For example, many non-Euclidean data are often represented as PSD matrices, e.g., brain connectivity analysis [35], diffusion tensor imaging [7, 24], and tomography [58]. In statistics, PSD matrices commonly appear as covariance matrices, typically assumed to be strictly positive definite (PD). In the analysis of such data or covariance estimation using the Riemannian geometry of the PD cone or its submanifolds, e.g., [8, 29, 72, 46, 51], the Euclidean metric is not suitable as geodesics leave the space in finite time, resulting in non-PSD interpolations. Therefore, various metrics have been proposed for the PD cone, including affine-invariant [60, 54, 48], log-Euclidean [3], log-Cholesky [40], Bures-Wasserstein [6, 65, 22], and a product metric with one metric on positive diagonal matrices and one metric on full-rank correlation matrices [66]. While these metrics are defined on the PD cone, the quotient geometry has been studied for fixed or bounded rank PSD matrices [67] and correlation matrices [15].
In covariance estimation based on the Riemannian frameworks, the choice of the parameter space and metric depends on the assumed covariance model and the data type. As an example, for $p$-dimensional vector data, the parameter space of the population covariance matrix $\Sigma$ is typically considered as the PD cone of the order $p$ [51, 29], denoted $\mathcal{S}_{p}^{++}$, where any aforementioned metric for the PD cone can be adopted. On the other hand, for $p_{1}\times p_{2}$ matrix-variate data, e.g., microarray data [2], phonetic data [56], and audio data [70], a separable (Kronecker) covariance model [18] is commonly used. Namely, for a zero-mean $p_{1}\times p_{2}$ random matrix $Y$, its $p_{1}p_{2}\times p_{1}p_{2}$ covariance matrix $\Sigma$ is formulated as
$\Sigma=V[Y]\equiv V[\text{vec}(Y)]=\Sigma_{2}\otimes\Sigma_{1},$ (1)
where $\Sigma_{1}\in\mathcal{S}_{p_{1}}^{++}$ and $\Sigma_{2}\in\mathcal{S}_{p_{2}}^{++}$ correspond to row and column covariance matrices, respectively. Here $\otimes$ denotes the Kronecker product. Note that we assume $p_{1},p_{2}\geq 2$ to emphasize the matrix structure of the data in this article. It follows that
@@
enabling the separate inference of correlation structures of row and column factors [18, 69]. This model is commonly used due to its parsimony and interpretability, involving at most $O(p_{1}^{2}+p_{2}^{2})$ correlations between variables. We denote the space of such separable covariance matrices by $\mathcal{S}_{p_{1},p_{2}}^{++}$. As a submanifold of $\mathcal{S}_{p}^{++}$ for $p:=p_{1}p_{2}$, [59, 46] proposed the estimation of the separable covariance matrix under the affine-invariant geometry.
@@
However, as $p$ grows, the separability assumption on $\Sigma$ as in (1) may oversimplify its correlation structure, allowing at most $O(p_{1}^{2}+p_{2}^{2})=o(p^{2})$ correlations, whereas $O(p^{2})$ correlations for the unstructured $\Sigma$. Hence, the separability assumption is often inappropriate, as also pointed out by [27, 26]. As a departure from this assumption, [33] introduced the core covariance matrix. They showed that every $\Sigma\in\mathcal{S}_{p}^{++}$ admits a unique decomposition into a separable component $K$, representing the most separable part of $\Sigma$, and a core component $C$, whitened $\Sigma$ via the identifiable square root $K^{1/2}$ of $K$, e.g., symmetric square root and Cholesky factor. Namely, $C=K^{-1/2}\Sigma K^{-1/2,\top}$ so that $\Sigma$ is represented as $K^{1/2}CK^{1/2,\top}$. This decomposition of $\Sigma$ is referred to as a Kronecker-core decomposition (KCD). Also, $\Sigma\in\mathcal{S}_{p_{1},p_{2}}^{++}$ if and only if $C=I_{p}$. Such a decomposition may also exist for rankr-r $\Sigma$ if $p_{1}/p_{2}+p_{2}/p_{1}<r$ [23, 19, 61], with $C$ sharing the same rank as $K^{1/2}$ is non-singular. By Proposition $5$ of [33], the dimension of the space of full-rank $C$ is $O(p^{2})$, whereas that of $\mathcal{S}_{p_{1},p_{2}}^{++}$ is $O(p_{1}^{2}+p_{2}^{2})=o(p^{2})$. Thus, in a high-dimensional regime where the sample size $n$ is smaller than the dimension of variables $p$, the estimation of $\Sigma$ is either numerically or statistically unstable without any structural assumption on $C$.
@@
As discussed by [33], one remedy is to introduce a partial-isotropy rankr-r structure to $C$, which commonly arises in factor analysis [5, 4]. Specifically, $\Omega\in\mathcal{S}_{p}^{++}$ has such a structure if $\lambda_{1}(\Omega)\geq\cdots\geq\lambda_{r}(\Omega)>\lambda_{r+1}(\Omega)=\cdots=\lambda_{p}(\Omega)>0$. Such $\Omega$ can be equivalently formulated as $AA^{\top}+cI_{p}$ for some $A\in\real^{p\times r}$ of full-column rank and constant $c>0$. Nevertheless, [33] did not pursue the estimation using this structure themselves, and left it as an open question, as characterizing such a core is crucial. In this article, we show that if $C$ exhibits a partial-isotropy rankr-r structure for fixed $r>p_{1}/p_{2}+p_{2}/p_{1}$, $C$ is a non-trivial convex combination of a rankr-r core and a trivial core $I_{p}$. The consequently proposed covariance model in this article is
$Y_{1},\ldots,Y_{n}\overset{i.i.d.}{\sim}N_{p_{1}\times p_{2}}(0,\Sigma)\text{ for }\Sigma=K^{1/2}((1-\lambda)AA^{\top}+\lambda I_{p})K^{1/2,\top},$ (2)
where $\lambda\in(0,1)$, the identifiable square root $K^{1/2}$ of $K$, and $A\in\real^{p\times r}$ of full-column rank such that $AA^{\top}$ is a core. We refer to the covariance model in (2) as a partial-isotropy core covariance model. As shown in Section 2.1 and 6, the coefficient $\lambda$ on $I_{p}$ quantifies how far $\Sigma$ is from being separable, improving the interpretability compared to an unstructured $C$.
As a remark, note that several two-way factor models have been proposed for matrix-variate data as a departure from the separable covariance model, e.g., Tucker factor model [13, 14, 16, 12] and canonical polyadic (CP) factor model [11, 30]. However, none of these models induce a natural measure of how far the true covariance is from being separable, since they do not continuously shrink toward the separable covariance model. To the best of our knowledge, the core shrinkage estimator proposed by [33] is only such an estimator via empirical Bayes. However, as demonstrated in Section 7, when the true core has a low-dimensional feature as above, this shrinkage estimator is subject to overparameterization. On the other hand, our proposed method directly exploits this feature and yields a natural measure via the estimate of the non-spiked eigenvalue of the true core.
To incorporate a partial-isotropy $C$ into the estimation of $\Sigma$ as in (2), we need a proper understanding of the space of rankr-r cores, denoted $\mathcal{C}_{p_{1},p_{2},r}^{+}$, motivating the study of its geometry. Therefore, this article is devoted to establishing the geometry of $\mathcal{C}_{p_{1},p_{2},r}^{+}$ and thus constructing the shrinkage estimator based on this geometry. Although ad hoc estimators may be used to estimate $\Sigma$ as in Section 7, we shall develop geometry to construct the estimator under the constraint that defines $C$. Specifically, we exploit the curvature of the negative log-likelihood to find the optimal $C$ that fits the data well.
@@
Our contributions are summarized in three main strands. First, we show that $\mathcal{C}_{p_{1},p_{2},r}^{+}$ is a compact, smooth, embedded submanifold of $\mathcal{S}_{p,r}^{+}$, the set of rankr-r PSD matrices. A key insight for the proof is that if $C=AA^{\top}\in\mathcal{C}_{p_{1},p_{2},r}^{+}$ for $A=[\text{vec}(A_{i}),\ldots,\text{vec}(A_{r})]$ with $A_{i}\in\real^{p_{1}\times p_{2}}$, Proposition $3$ of [33] implies that
@@
Given $p_{1}/p_{2}+p_{2}/p_{1}<r$, we construct $\mathcal{C}_{p_{1},p_{2},r}^{+}$ as the smooth image of the smooth manifold $\mathcal{D}_{p_{1},p_{2},r}$ consisting of $\tilde{A}=(A_{1},\ldots,A_{r})$ satisfying the above. While the proof is straightforward when $r=p$, the rank-deficient case requires additional technical work. Namely, the canonically decomposable $\tilde{A}$, i.e., there exist non-singular matrices $(P,Q)$ such that $PA_{i}Q^{-1}$ is of non-trivial block-diagonal form, prevents $\mathcal{D}_{p_{1},p_{2},r}$ from being a smooth manifold. The canonical decomposability notion arises in the study of the threshold on $r$ for which a generic $\Omega\in\mathcal{S}_{p,r}^{+}$ admits the Kronecker MLE [19, 61] and hence the KCD. To say informally, the canonically indecomposable $\tilde{A}$ guarantees that the row factors and column factors are well-connected, which is made more precise in Section 4.1. We illustrate in Section 4.1 with an example how the set of canonically decomposable matrices prevents $\mathcal{D}_{p_{1},p_{2},r}$ from being a smooth manifold and is a closed set of measure zero, motivating its removal. The proof strategy is first outlined with the case $r=p$, where $\mathcal{C}_{p_{1},p_{2}}^{++}\equiv\mathcal{C}_{p_{1},p_{2},p}^{+}$, in Section 3, and then extended to the case where $r<p$ in Section 4.2.
@@
The next contribution is to study the differential geometry of $\mathcal{C}_{p_{1},p_{2}}^{++}$, $\mathcal{C}_{p_{1},p_{2},r}$, and the quotient manifold $\mathcal{C}_{p_{1},p_{2},r}/\mathcal{O}_{r}$, which also serves as the ingredients for manifold optimization to compute the covariance estimator incorporating a partial-isotropy core. Let $k(\Sigma)$ and $c(\Sigma)$ denote the separable and core components of $\Sigma$, respectively. We refer to $k$ and $c$ as the Kronecker and core maps defined on $\mathcal{S}_{p}^{++}$, respectively. With the map $f:\Sigma\in\mathcal{S}_{p}^{++}\rightarrow(k(\Sigma),c(\Sigma))\in\mathcal{S}_{p_{1},p_{2}}^{++}\times\mathcal{C}_{p_{1},p_{2}}^{++}$, we show that $\mathcal{S}_{p}^{++}$ is diffeomorphic to $\mathcal{S}_{p_{1},p_{2}}^{++}\times\mathcal{C}_{p_{1},p_{2}}^{++}$ via the map $f$ in Section 5.1. Therefore, we provide a new insight into the smooth structure of $\mathcal{S}_{p}^{++}$ via separability. We also compute the differentials of $f$ and its inverse $f^{-1}$. Under the Euclidean metric, we derive the Riemannian gradient and Hessian operator on $\mathcal{C}_{p_{1},p_{2}}^{++}$ in Section 5.2. The same is done for $\mathcal{C}_{p_{1},p_{2},r}$ under the same metric, which we employ in manifold optimization, and for the quotient manifold $\mathcal{C}_{p_{1},p_{2},r}/\mathcal{O}_{r}$ under the induced quotient metric in Section 5.3.
@@
Finally, using the geometry of $\mathcal{C}_{p_{1},p_{2},r}$, we propose a partial-isotropy core shrinkage estimator (PICSE) in Section 6, assuming the covariance model in (2) on the data. This answers the open question posed by [33] on how a partial-isotropy core can be incorporated into estimating $\Sigma$. We provide an alternating minimization procedure of the negative log-likelihood in the parameters $(K^{1/2},A,\lambda)$ given in (2) to compute PICSE. In updating $A$, we leverage the curvature of the objective function on $\mathcal{C}_{p_{1},p_{2},r}$ via second-order Riemannian manifold optimization using the results in Section 5.3, with some suitable retraction. In Section 7, we numerically illustrate that PICSE outperforms existing covariance estimators for matrix-variate data, and some baseline methods.
@@
The rest of the article is organized as follows. Section 1.1 introduces notations used throughout this article. In Section 2, we review some preliminaries, including the KCD, Riemannian manifolds, quotient manifolds, and algebraic geometry. In Section 3, we prove that $\mathcal{C}_{p_{1},p_{2}}^{++}$ is a compact, smooth, embedded submanifold of $\mathcal{S}_{p}^{++}$. When $p_{1}/p_{2}+p_{2}/p_{1}<r<p$, it is shown that $\mathcal{C}_{p_{1},p_{2},r}^{+}$ is a compact, smooth, embedded submanifold of $\mathcal{S}_{p,r}^{+}$ in Section 4 after removing the set of canonically decomposable matrices, using the proof strategy developed in Section 3. In Section 5, we establish the diffeomorphic relationship between $\mathcal{S}_{p}^{++}$ and $\mathcal{S}_{p_{1},p_{2}}^{++}\times\mathcal{C}_{p_{1},p_{2}}^{++}$. We also derive the differential geometric quantities relevant to $\mathcal{C}_{p_{1},p_{2}}^{++}$, $\mathcal{C}_{p_{1},p_{2},r}$, and $\mathcal{C}_{p_{1},p_{2},r}/\mathcal{O}_{r}$ under the Euclidean metric. Leveraging the geometry of $\mathcal{C}_{p_{1},p_{2},r}$, the partial isotropy core shrinkage estimator (PICSE) is proposed in Section 6, supported by numerical illustrations in Section 7. Section 8 concludes the article with a discussion. All the omitted proofs are deferred to Appendix A. The formulas of Euclidean derivative and Hessian operator associated with computing PICSE are provided in Appendix B. The additional figures and tables that illustrate the results in Section 7 are given in Appendix C.
1.1 Notations